QUESTION 36

Table 4.2. Production Costs and Prices of Imported and Domestic TVs

Imported TVs

Domestic TVs

Component parts $600 Imported component parts $600
Assembly cost/profit    150 Assembly cost 150
Nominal tariff     50 Profit     50
Import price Domestic price
after tariff 800 after tariff 800

Consider Table 4.2. Prior to the tariff, the total price of imported TVs is:

$750

$600

$300

$700

QUESTION 38

  1. Table 4.1. Production Costs and Prices of Imported and Domestic VCRs

    Imported VCRs

    Domestic VCRs

    Component parts $150 Imported component parts $150
    Assembly cost/profit     50 Assembly cost     50
    Nominal tariff     25 Profit     25
    Import price Domestic price
    after tariff 225 after tariff 225

    Consider Table 4.1. Prior to the tariff, the total price of imported VCRs is:

    $150

    $200

    $225

    $235
Answer #1

1.Total price without tariff=Component parts+assembly cost

Total price=600+150=750

Answer-750

2.Total price of imported VCR before tariff=Component parts+assembly cost

Total price=150+50=200

Answer-200